# Alternating geometric progressions modulo one and Sturmian words

### 23 Mar 2026 v1

@@ 1 Introduction
Distribution modulo one of geometric progressions has been the subject of extensive studies. The famous question considered by [Mahler] on the case of common ratio $3/2$ remains open. A theorem of Dubickas [Dubickas-BLMS, Theorem 1] implies that for any integer $b\geq 2$ and any irrational real number $\xi$, the numbers $\{\xi b^{n}\}$, $n\geq 0$ cannot be contained in an interval of length less than $1/b$. Here $\{x\}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}x-\lfloor x\rfloor$ denotes the fractional part of a real number $x$. Bugeaud and Dubickas [BD, Theorem 2.1] showed that for an irrational real number $\xi$, the numbers $\{\xi b^{n}\}$, $n\geq 0$ are contained in an interval of length $1/b$ if and only if the representation $w_{1}w_{2}\dots$ of $\{\xi\}$ in base $b$ is given by a Sturmian word $w_{1}w_{2}\dots$ over an alphabet consisting of two adjacent integers. As pointed out in [AG], the case $b=2$ was proved by Veerman [Veerman, Theorem 2.1] and rediscovered several times.
For alternating geometric progressions with integral common ratio, we showed in previous work [LZ, Theorem 1.1] that for any irrational real number $\xi$, $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is not contained in any interval of $\mathbb{R}/\mathbb{Z}$ of length less than $b^{-1}+b^{-2}-b^{-3}$. Here $\pi\colon\mathbb{R}\to\mathbb{R}$$\mathbb{Z}$ denotes the projection. In this paper, we describe the irrational real numbers $\xi$ such that $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval of $\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}$.
To state our results we need some terminology from combinatorics on words. An infinite word $w_{1}w_{2}\dots$ is said to be aperiodic if there does not exist integers $k,N\geq 1$ such that $w_{n}=w_{n+k}$ for all $n\geq N$. The slope of an integer-valued infinite word $w_{1}w_{2}\dots$ is defined to be $\lim_{n\to\infty}\frac{\sum_{i=1}^{n}w_{i}}{n}$. An integer-valued infinite word with irrational slope is necessarily aperiodic. An infinite word $s_{1}s_{2}\dots$ over the alphabet $\{0,1\}$ is Sturmian if and only if there exists an irrational $\theta\in(0,1)$ and a $\rho\in\mathbb{R}$ such that
@@
In negative-base number systems, the order on real numbers is compatible with the alternate order $\preceq_{\mathrm{alt}}$ on words rather than the lexicographical order $\preceq_{\mathrm{lex}}$. For integer-valued infinite words $\mathbf{v}=v_{1}v_{2}\dots$ and $\mathbf{w}=w_{1}w_{2}\dots$, we write $\mathbf{v}\preceq_{\mathrm{alt}}\mathbf{w}$ if $N\mathbf{v}\preceq_{\mathrm{lex}}N\mathbf{w}$, where $N\mathbf{v}=(-v_{1})v_{2}(-v_{3})v_{4}\dots$ and similarly for $N\mathbf{w}$. It is easy to see that, for infinite words $\mathbf{v}$ and $\mathbf{w}$ over $\{0,1\}$ and $0<r<1/2$, $\mathbf{v}\preceq_{\mathrm{alt}}\mathbf{w}$ if and only if $t_{-r}(\mathbf{v})\leq t_{-r}(\mathbf{w})$ (Lemma 4.7). Theorem 1.2 implies the following characterization of the class $\mathcal{D}$.
@@
For alternating geometric progressions with rational common ratio p/q-p/q, where $p>q>1$ are integers, we showed in [LZ, Theorem 1.1] that for any real number $\xi\neq 0$, the sequence $(\pi(\xi(-p/q)^{n}))_{n\geq 0}$ is not contained in any interval of $\mathbb{R}/\mathbb{Z}$ of length less than $(r+r^{2}-r^{3})/q$. In subsequent work [LZ3], we will use Theorem 1.4 to study the question whether the sequence can lie in an interval of length equal to $(r+r^{2}-r^{3})/q$. The analogous question for powers of positive rational numbers was studied by several authors ([FLP], [Bugeaud-linear], [Dubickas-small], [Dubickas-large]).
@@ 3 The positive case
By Lemma 2.6, $t_{r}(0\mathbf{c}_{\theta})\in W$ if and only if $T^{n}\mathbf{w}=0\mathbf{c}_{\theta}$ for some $n\geq 0$, and $t_{r}(1\mathbf{c}_{\theta})\in W$ if and only if $T^{m}\mathbf{w}=1\mathbf{c}_{\theta}$ for some $m\geq 0$. If both conditions are satisfied, then $T^{m+1}\mathbf{w}=\mathbf{c}_{\theta}=T^{n+1}\mathbf{w}$ for $m\neq n$, which contradicts the aperiodicity of $\mathbf{w}$. Thus $W$ is contained in $t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta}))$ or $(t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta})]$. Furthermore, $W$ is contained in $(t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta}))$ if and only if $T^{n}\mathbf{w}=\mathbf{c}_{\theta}$ for some $n\geq 1$. ∎
@@
Here $\theta$ denotes the slope of $\mathbf{w}$, and $\inf^{\mathrm{lex}}$ and $\sup^{\mathrm{lex}}$ denote respectively the infimum and supremum with respect to the order $\preceq_{\mathrm{alt}}$.
@@
This is essentially [BD, Theorem 2.1], except that here we do not a priori exclude the case where $\pi(0)$ is in the interior of the interval or is the upper endpoint. Assume that $(\pi(\xi b^{n}))_{n\geq 0}$ is contained in an interval $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $1/b$. Choose $\eta\in\mathbb{R}$ such that $\pi(\eta)\notin I$. Then $I$ lifts to a unique interval $\tilde{I}\subseteq(\eta,\eta+1)$. For $n\geq 0$, let $y_{n}=\{\xi b^{n}-\eta\}+\eta\in(\eta,\eta+1)$ and $v_{n}=by_{n}-y_{n+1}\in\mathbb{Z}\cap((b-1)\eta-1,(b-1)\eta+b)$. Since $\pi(y_{n})=\pi(\xi b^{n})$, we have $y_{n}\in\tilde{I}$ for all $n\geq 0$. Let $\mathbf{v}=v_{0}v_{1}\dots$. Then $t_{1/b}(T^{n}\mathbf{v})=\sum_{i=0}^{\infty}v_{n+i}b^{-(i+1)}=y_{n}$. Since $y_{0}$ is irrational, $\mathbf{v}$ is aperiodic. Then, by Proposition 3.1, there exists $g\in\mathbb{Z}$ and a Sturmian word $\mathbf{w}=w_{0}w_{1}\dots$ over $\{0,1\}$ such that $v_{n}=g+w_{n}$ for all $n\geq 0$. Thus $\pi(\xi)=\pi(y_{0})=\pi(g/(b-1)+t_{1/b}(\mathbf{w}))$. By a theorem of Ferenczi and Mauduit [FM, Proposition 2], $t_{1/b}(\mathbf{w})$ is transcendental and consequently the same holds for $\xi$. The other assertions of Corollary 3.3 follow easily from the corresponding assertions of Proposition 3.1 and the formula $\pi(t_{1/b}(T^{n}\mathbf{v}))=\pi(b^{n}t_{1/b}(\mathbf{v}))$. ∎
@@ 4 Proofs of the main results
Let $\mathbf{w}$ be an aperiodic word over $\{0,1\}$. Then $\mathbf{w}\in\mathcal{S}$ if and only there does not exist a finite word $\mathbf{u}$ such that $01\mathbf{u}1$ and $10\mathbf{u}0$ are subwords of $\mathbf{w}$.
@@
$t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})=(t_{r}(\mathbf{u}^{\prime})-t_{r}(\mathbf{u}))+r^{n}(t_{r}(T^{n}\mathbf{s}^{\prime})-t_{r}(T^{n}\mathbf{s}))$
@@
and $t_{r^{2}}(\mathbf{u})\in[t_{r^{2}}(0^{\infty}),t_{r^{2}}(1^{\infty})]$, where $a\in\Sigma$ and $\mathbf{u}\in\Sigma^{\mathbb{N}}$. The equality conditions follow from the fact that $t_{r^{2}}(\mathbf{u})=t_{r^{2}}(a^{\infty})$ if and only if $\mathbf{u}=a^{\infty}$. ∎
@@
We apply Theorem 1.4 to $r=1/b$ as follows. Assume that $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}<1$. Choose $\eta\in\mathbb{R}$ such that $\pi(\eta)\notin I$. Then $I$ lifts to a unique interval $\tilde{I}\subseteq(\eta,\eta+1)$. For $n\geq 0$, let $y_{n}=\{\xi(-b)^{n}-\eta\}+\eta\in(\eta,\eta+1)$ and $v_{n}=-by_{n}-y_{n+1}\in\mathbb{Z}\cap(-(b+1)(\eta+1),-(b+1)\eta)$. Since $\pi(y_{n})=\pi(\xi(-b)^{n})$, we have $y_{n}\in\tilde{I}$ for all $n\geq 0$. Let $\mathbf{v}=v_{0}v_{1}\dots$. Then $t_{-1/b}(T^{n}\mathbf{v})=\sum_{i=0}^{\infty}v_{n+i}(-b)^{-(i+1)}=y_{n}$. Since $y_{0}$ is irrational, $\mathbf{v}$ is aperiodic. Then, by Theorem 1.4, there exist $g\in\mathbb{Z}$ and $\mathbf{w}=w_{1}w_{2}\dotso\in\mathcal{D}$ such that $v_{i}=-g+w_{i}$ for all $i\geq 1$. Thus $\pi(\xi)=\pi(y_{0})=\pi(g/(b+1)+t_{-1/b}(\mathbf{w}))$. By Lemma 4.11, it follows that $\xi$ is transcendental. The other assertions of Theorem 1.2 follow easily from the corresponding assertions of Theorem 1.4 and the formula $\pi(t_{-1/b}(T^{n}\mathbf{v}))=\pi((-b)^{n}t_{-1/b}(\mathbf{v}))$. ∎

### 02 Apr 2026 v2

@@ 1 Introduction
Distribution modulo one of geometric progressions has been the subject of extensive studies. The famous question considered by Mahler [Mahler] concerning the case of common ratio $3/2$ remains open. A theorem of Dubickas [Dubickas-BLMS, Theorem 1] implies that for any integer $b\geq 2$ and any irrational real number $\xi$, the numbers $\{\xi b^{n}\}$, $n\geq 0$ cannot be contained in an interval of length less than $1/b$. Here $\{x\}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}x-\lfloor x\rfloor$ denotes the fractional part of a real number $x$. Bugeaud and Dubickas [BD, Theorem 2.1] showed that for an irrational real number $\xi$, the numbers $\{\xi b^{n}\}$, $n\geq 0$ are contained in an interval of length equal to $1/b$ if and only if the representation $w_{1}w_{2}\dots$ of $\{\xi\}$ in base $b$ is given by a Sturmian word $w_{1}w_{2}\dots$ over an alphabet consisting of two adjacent integers. As pointed out in [AG], the case $b=2$ was proved by Veerman [Veerman, Theorem 2.1] and rediscovered several times.
For alternating geometric progressions with integral common ratio, we showed in previous work [LZ, Theorem 1.1] that for any irrational real number $\xi$, $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is not contained in any interval of $\mathbb{R}/\mathbb{Z}$ of length less than $b^{-1}+b^{-2}-b^{-3}$. Here $\pi\colon\mathbb{R}\to\mathbb{R}$$\mathbb{Z}$ denotes the projection. In this paper, we describe the irrational real numbers $\xi$ such that $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval of $\mathbb{R}/\mathbb{Z}$ of length equal to $b^{-1}+b^{-2}-b^{-3}$.
To state our results we need some terminology from combinatorics on words. An infinite word $w_{1}w_{2}\dots$ is said to be aperiodic if there does not exist integers $k,N\geq 1$ such that $w_{n}=w_{n+k}$ for all $n\geq N$. The slope of an infinite word $w_{1}w_{2}\dots$ over the alphabet $\mathbb{Z}$ is defined to be $\lim_{n\to\infty}\frac{\sum_{i=1}^{n}w_{i}}{n}$. An infinite word over $\mathbb{Z}$ with irrational slope is necessarily aperiodic. An infinite word $s_{1}s_{2}\dots$ over the alphabet $\{0,1\}$ is Sturmian if and only if there exists an irrational $\theta\in(0,1)$ and a $\rho\in\mathbb{R}$ such that
@@
In negative-base number systems, the order on real numbers is compatible with the alternate order $\preceq_{\mathrm{alt}}$ on words rather than the lexicographical order $\preceq_{\mathrm{lex}}$. For infinite words $\mathbf{v}=v_{1}v_{2}\dots$ and $\mathbf{w}=w_{1}w_{2}\dots$ over $\mathbb{Z}$, we write $\mathbf{v}\preceq_{\mathrm{alt}}\mathbf{w}$ if $N\mathbf{v}\preceq_{\mathrm{lex}}N\mathbf{w}$, where $N\mathbf{v}=(-v_{1})v_{2}(-v_{3})v_{4}\dots$ and similarly for $N\mathbf{w}$. It is easy to see that, for infinite words $\mathbf{v}$ and $\mathbf{w}$ over $\{0,1\}$ and $0<r<1/2$, $\mathbf{v}\preceq_{\mathrm{alt}}\mathbf{w}$ if and only if $t_{-r}(\mathbf{v})\leq t_{-r}(\mathbf{w})$ (Lemma 4.7). Theorem 1.2 implies the following characterization of the class $\mathcal{D}$.
@@
For alternating geometric progressions with rational common ratio p/q-p/q, where $p>q>1$ are integers, we showed in [LZ, Theorem 1.1] that for any real number $\xi\neq 0$, the sequence $(\pi(\xi(-p/q)^{n}))_{n\geq 0}$ is not contained in any interval of $\mathbb{R}/\mathbb{Z}$ of length less than $(r+r^{2}-r^{3})/q$, where $r=q/p$. In subsequent work [LZ3], we will use Theorem 1.4 to study the question whether the sequence can lie in an interval of length equal to $(r+r^{2}-r^{3})/q$. The analogous question for powers of positive rational numbers was studied by several authors ([FLP], [Bugeaud-linear], [Dubickas-small], [Dubickas-large]).
@@ 3 The positive case
By Lemma 2.6, $t_{r}(0\mathbf{c}_{\theta})\in W$ if and only if $T^{n}\mathbf{w}=0\mathbf{c}_{\theta}$ for some $n\geq 0$, and $t_{r}(1\mathbf{c}_{\theta})\in W$ if and only if $T^{m}\mathbf{w}=1\mathbf{c}_{\theta}$ for some $m\geq 0$. If both conditions are satisfied, then $T^{m+1}\mathbf{w}=\mathbf{c}_{\theta}=T^{n+1}\mathbf{w}$ for $m\neq n$, which contradicts the aperiodicity of $\mathbf{w}$. Thus $W$ is contained in $t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta}))$ or $(t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta})]$. Furthermore, $W$ is not contained in $(t_{r}(0\mathbf{c}_{\theta}),t_{r}(1\mathbf{c}_{\theta}))$ if and only if $T^{n}\mathbf{w}=\mathbf{c}_{\theta}$ for some $n\geq 1$. ∎
@@
Here $\theta$ denotes the slope of $\mathbf{w}$, and $\inf^{\mathrm{lex}}$ and $\sup^{\mathrm{lex}}$ denote respectively the infimum and supremum with respect to the order $\preceq_{\mathrm{lex}}$.
@@
This is essentially [BD, Theorem 2.1], except that here we do not a priori exclude the case where $\pi(0)$ is in the interior of the interval or is the upper endpoint. Assume that $(\pi(\xi b^{n}))_{n\geq 0}$ is contained in an interval $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $1/b$. Choose $\eta\in\mathbb{R}$ such that $\pi(\eta)\notin I$. Then $I$ lifts to a unique interval $\tilde{I}\subseteq(\eta,\eta+1)$ of length $1/b$. For $n\geq 0$, let $y_{n}=\{\xi b^{n}-\eta\}+\eta\in(\eta,\eta+1)$ and $v_{n}=by_{n}-y_{n+1}\in\mathbb{Z}\cap((b-1)\eta-1,(b-1)\eta+b)$. Since $\pi(y_{n})=\pi(\xi b^{n})$, we have $y_{n}\in\tilde{I}$ for all $n\geq 0$. Let $\mathbf{v}=v_{0}v_{1}\dots$. Then $t_{1/b}(T^{n}\mathbf{v})=\sum_{i=0}^{\infty}v_{n+i}b^{-(i+1)}=y_{n}$. Since $y_{0}$ is irrational, $\mathbf{v}$ is aperiodic. Then, by Proposition 3.1, there exists $g\in\mathbb{Z}$ and a Sturmian word $\mathbf{w}=w_{0}w_{1}\dots$ over $\{0,1\}$ such that $v_{n}=g+w_{n}$ for all $n\geq 0$. Thus $\pi(\xi)=\pi(y_{0})=\pi(g/(b-1)+t_{1/b}(\mathbf{w}))$. By a theorem of Ferenczi and Mauduit [FM, Proposition 2], $t_{1/b}(\mathbf{w})$ is transcendental and consequently the same holds for $\xi$. The other assertions of Corollary 3.3 follow easily from the corresponding assertions of Proposition 3.1 and the formula $\pi(t_{1/b}(T^{n}\mathbf{v}))=\pi(b^{n}t_{1/b}(\mathbf{v}))$. ∎
@@ 4 Proofs of the main results
Let $\mathbf{w}$ be an aperiodic word over $\{0,1\}$. Then $\mathbf{w}\in\mathcal{S}$ if and only if there does not exist a finite word $\mathbf{u}$ such that $01\mathbf{u}1$ and $10\mathbf{u}0$ are subwords of $\mathbf{w}$.
@@
$t_{r}(\mathbf{s}^{\prime})-t_{r}(\mathbf{s})=(t_{r}(\mathbf{u}^{\prime})-t_{r}(\mathbf{u}))+r^{n}(t_{r}(T^{n}\mathbf{s}^{\prime})-t_{r}(T^{n}\mathbf{s})).$
@@
and $t_{r^{2}}(\mathbf{u})\in[t_{r^{2}}(0^{\infty}),t_{r^{2}}(1^{\infty})]$, where $a\in\Sigma$ and $\mathbf{u}\in\Sigma^{\mathbb{N}}$. The equality conditions follow from the fact that, for $a\in\{0,1\}$, $t_{r^{2}}(\mathbf{u})=t_{r^{2}}(a^{\infty})$ if and only if $\mathbf{u}=a^{\infty}$. ∎
@@
We apply Theorem 1.4 to $r=1/b$ as follows. Assume that $(\pi(\xi(-b)^{n}))_{n\geq 0}$ is contained in an interval $I\subseteq\mathbb{R}/\mathbb{Z}$ of length $b^{-1}+b^{-2}-b^{-3}<1$. Choose $\eta\in\mathbb{R}$ such that $\pi(\eta)\notin I$. Then $I$ lifts to a unique interval $\tilde{I}\subseteq(\eta,\eta+1)$ of length $b^{-1}+b^{-2}-b^{-3}$. For $n\geq 0$, let $y_{n}=\{\xi(-b)^{n}-\eta\}+\eta\in(\eta,\eta+1)$ and $v_{n}=-by_{n}-y_{n+1}\in\mathbb{Z}\cap(-(b+1)(\eta+1),-(b+1)\eta)$. Since $\pi(y_{n})=\pi(\xi(-b)^{n})$, we have $y_{n}\in\tilde{I}$ for all $n\geq 0$. Let $\mathbf{v}=v_{0}v_{1}\dots$. Then $t_{-1/b}(T^{n}\mathbf{v})=\sum_{i=0}^{\infty}v_{n+i}(-b)^{-(i+1)}=y_{n}$. Since $y_{0}$ is irrational, $\mathbf{v}$ is aperiodic. Then, by Theorem 1.4, there exist $g\in\mathbb{Z}$ and $\mathbf{w}=w_{0}w_{1}\dotso\in\mathcal{D}$ such that $v_{i}=-g+w_{i}$ for all $i\geq 0$. Thus $\pi(\xi)=\pi(y_{0})=\pi(g/(b+1)+t_{-1/b}(\mathbf{w}))$. By Lemma 4.11, it follows that $\xi$ is transcendental. The other assertions of Theorem 1.2 follow easily from the corresponding assertions of Theorem 1.4 and the formula $\pi(t_{-1/b}(T^{n}\mathbf{v}))=\pi((-b)^{n}t_{-1/b}(\mathbf{v}))$. ∎